OpenStudy (anonymous):

PLEASE HELP WILL GIVE A MEDAL Find the derivative of f(x) = 4/x at x = 2.

OpenStudy (anonymous):

@zepdrix

Mehek (mehek14):

so replace x with 2 f(x) = 4/2 = 2

OpenStudy (freckles):

are we suppose to use the definition or can use product or quotient rule?

OpenStudy (freckles):

or you know like the short cut rules

OpenStudy (freckles):

like I would just use power rule here

OpenStudy (freckles):

\[f(x)=4x^{-1} \\ f'(x)=4 \cdot ?\]

OpenStudy (freckles):

I will let you do the power rule thing :)

OpenStudy (anonymous):

im confused how do I apply that?

OpenStudy (freckles):

if you have y=x^n then y'=nx^(n-1)

OpenStudy (freckles):

the n here is -1

OpenStudy (anonymous):

so 4/2^2?

OpenStudy (freckles):

well almost

OpenStudy (freckles):

I think you forgot to bring down the -1

OpenStudy (freckles):

\[f(x)=4x^{-1} \\ f'(x)=4(-1)x^{-1-1}=-4x^{-2}=\frac{-4}{x^2} \]

OpenStudy (anonymous):

ohhh so it would be -1?

OpenStudy (freckles):

the power is -1

OpenStudy (freckles):

so yeah n is -1

OpenStudy (anonymous):

do you mind helping me with another one similar to this one?